# Nigel - Mechatronic Design and Robust Sim2Real Control of an Over-Actuated Autonomous Vehicle

### v1

@@ Abstract
Simulation to reality (sim2real) transfer from a dynamics and controls perspective usually involves re-tuning or adapting the designed algorithms to suit real-world operating conditions, which often violates the performance guarantees established originally. This work presents a generalizable framework for achieving reliable sim2real transfer of autonomy-oriented control systems using multi-model multi-objective robust optimal control synthesis, which lends well to uncertainty handling and disturbance rejection with theoretical guarantees. Particularly, this work is centered around an actuation-redundant scaled autonomous vehicle called Nigel, with independent all-wheel drive and independent all-wheel steering architecture, whose enhanced configuration space bodes well for robust control applications. To this end, we present a systematic study on the complete mechatronic design, dynamics modeling, parameter identification, and robust stabilizing as well as steady-state tracking control of Nigel using the proposed framework, with experimental validation.
@@
Autonomous vehicles, 4WD4WS vehicles, over-actuated systems, mechatronic design, robust optimal control, linear matrix inequalities, sim2real transfer, uncertainty handling, disturbance rejection.
@@ I Introduction
Figure 1: Nigel: A mechatronically redundant and reconfigurable 1:14 scale autonomous vehicle. This modular and open-source vehicle platform offers multiple driving/steering configurations, abundant interoceptive/exteroceptive sensors, on-board edge-AI computation and wired/wireless communication, selectively constrained software-defined actuation, as well as a fully functional lighting and signaling system. A video highlighting key features of Nigel is available at https://youtu.be/UVIShZuZmpg
Modern day automotive systems exploit a combination of mechanical electrical, electronic, networking, and software sub-systems to enhance performance via a hierarchical suite of autonomy-oriented control realizations. While earlier autonomy developers may have enjoyed the freedom of primarily focusing on core software development, the present context demands a paradigm shift towards a synergistic hardware-software co-design approach, harmonizing with the principles of mechatronics engineering [1]. In particular, the demand for increased maneuverability, enhanced control configuration space, and improved tolerance against faults motivates the pursuit of unconventional vehicle designs. In response, autonomous vehicle (AV) realizations must augment their underlying hardware-software architectures to enhance core performance as well as adaptability to changing operating conditions. However, with the advent of novel design architectures, advanced control strategies [2] are required to fully exploit the added capabilities. Moreover, the devised control systems need to successfully transition the sim2real gap by guaranteeing robust performance against real-world uncertainties and disturbances. This further calls for model and simulation-based controller refinement along with real-world testing, which requires capable cyber-physical deployment platforms of varying scales. Previous works have tried to address some of these aspects, albeit in a fragmented sense as we will review below.
@@
Platform ${}^{\mathrm{\dagger}}$ Sensing Modalities Computational Resources
@@
Monocular Camera
Depth/Stereo Camera
@@
MIT Racecar 1:10 \$2,600 Jetson TX2 VESC AS
AutoRally 1:5 \$23,300 Custom Teensy LC/Arduino Micro AS
F1TENTH 1:10 \$3,260 Jetson TX2 VESC 6MkV AS
DSV 1:10 \$1,000 ODROID-XU4 Arduino (Mega + Uno) AS
MuSHR 1:10 \$930 Jetson Nano Turnigy SK8-ESC AS
BARC 1:10 \$1,030 ODROID-XU4 Arduino Nano AS
ORCA 1:43 \$960 None ARM Cortex M4 $\mu$C AS
HyphaROS 1:10 \$600 ODROID-XU4 RC ESC TBLE-02S AS
Donkey Car 1:16 \$370 Raspberry Pi ESC AS
QCar 1:10 \$20,000 Jetson TX2 Proprietary AS
DeepRacer 1:18 \$400 Proprietary Proprietary AS
Duckiebot N/A \$450 Raspberry Pi/Jetson Nano None DD
TurtleBot3 N/A \$590 Raspberry Pi OpenCR DD
${}^{\mathrm{\dagger}}$ll cost values are ceiled to the nearest \$10. ${}^{\mathrm{\ddagger}}$ctuation mechanisms comprise Ackermann steered (AS), differential-drive (DD), and 4-wheel drive 4-wheel steer (4WD4WS) configurations.
@@
In the context of developing scaled autonomous vehicles, educational and research institutes have contributed platforms like the MIT Racecar [3], AutoRally [4], F1TENTH [5], Multi-agent System for non-Holonomic Racing (MuSHR) [6], Optimal RC Racing (ORCA) Project [7], Delft Scaled Vehicle (DSV) [8] and Berkeley Autonomous Race Car (BARC) [9]. Community-driven projects such as HyphaROS RaceCar [10] and Donkey Car [11] have also emerged, which are often tailored for specific applications like map-based navigation and vision-aided imitation learning, respectively. Moreover, commercial products like Quanser QCar [12] and AWS DeepRacer [13] have also entered the market, but their closed-source nature and prohibitive costs limit their accessibility to the broader community. Additionally, scaled robot platforms like the TurtleBot3 [14] and Duckiebot [15] remain valuable for teaching fundamental autonomy concepts. However, none of the aforementioned platforms (refer Table I) contributes primarily towards novel vehicular configuration or architecture. Other works such as [16] and [17] have recently prototyped over-actuated scaled vehicles with the primary aim of validating their control algorithms, and as such, none of these platforms are open-sourced. Additionally, both of these platforms lack comprehensive autonomy features and only the latter offers truly independent driving and steering capability, although at a much larger scale (1:5) and without extended steering limits. To the best of authors’ knowledge, Nigel (refer Fig. 1) is the first open-source [^1] GitHub: https://github.com/AutoDRIVE-Ecosystem mechatronically redundant [18] autonomous vehicle platform offering comprehensive autonomy features as well as independent all-wheel driving and independent all-wheel steering (i.e., independent 4WD4WS) configuration with extended (±$\pm 90^{\circ}$) steering angles, within a small footprint of 1:14 scale. Additionally, Nigel is a part of the larger AutoDRIVE Ecosystem [^2] Website: https://autodrive-ecosystem.github.io [19, 20], which also offers a high-fidelity digital-twin simulation platform [21, 22], as well as flexible application programming interfaces (APIs) to develop low and high-level autonomy algorithms.
@@
In terms of developing control strategies for vehicles with unconventional architectures, prior research endeavors have proposed and applied techniques such as inverse dynamics control [23], coordinated motion control [24], adaptive steering control [25] and model predictive control [26, 27] for achieving driver-assistance as well as automatic control objectives. Although some of these works discuss the adaptability and robustness of the designed controllers in simulated experiments, they cannot guarantee similar performance in real-world conditions. Other recent works such as [28] and [29] have applied cascaded feedforward control and robust $H_{\infty}$ control, respectively, to handle environmental uncertainties and disturbances using over-actuated vehicle architectures. These robust control techniques in themselves are not completely novel, and have been studied and applied to automotive systems previously for yaw-plane [30, 31, 32], roll-plane [33] or vertical dynamics stabilization [34], automatic emergency braking [35], trajectory tracking [36], etc.
@@
Our work poses the problem of robust control from a “sim2real transfer” lens, to develop a generalizable framework for reliably bridging the dynamics interface of the sim2real gap. To this end, we demonstrate robust stabilizing as well as steady-state tracking control of Nigel with seamless sim2real transfer, by designing a multi-model multi-objective full-state feedback controller using a linear matrix inequality (LMI) approach, which seeks the optimal tradeoff between $H_{2}$ and $H_{\infty}$ performance with $D$-stability. It is to be noted that, contrary to some of our previous works, which have targeted bridging the sim2real gap by increasing simulation fidelity [37], this work intentionally widens the reality gap by using a simplified linear dynamics model with lumped wheel representation to synthesize the robust optimal controller, which is hypothesized to bridge the dynamics gap during real-world deployment. As such, the devised controller is tested rigorously in simulation as well as real-world settings for different benchmark maneuvers under varying grades of disturbances and uncertainties. In addition to validating the sim2real transfer, we also analyze the controller performance against deliberately injected exogenous disturbances and uncertainties during real-world experiments. The key contributions of this paper are summarized below:
@@
  • Open-hardware, open-software design architecture of a novel 1:14 scale independent 4WD4WS autonomous vehicle, with extended steering limits is presented.
@@
  • Non-linear dynamics model of such a 4WD4WS vehicle is derived and transformed into a linear uncertain system model with reasonable parameter identification.
@@
  • Robust sim2real control framework is established and validated using exhaustive experiments in simulation as well as real-world settings.
The remainder of this paper is organized as follows. Section II elucidates the complete mechatronic design architecture of Nigel. A detailed derivation of the independent 4WD4WS vehicle dynamics model and its analysis is discussed in Section III. Formulation of the proposed generalizable robust optimal sim2real control framework is presented in Section IV. Details pertaining to its performance evaluation in simulation as well as real-world settings are discussed in Section V. Finally, Section VI presents concluding remarks and future directions.
@@ II Vehicle Design Architecture
Figure 2: Various components and sub-systems of Nigel. Dashed lines indicate occluded or hidden components.
@@
Chassis: Nigel is a 1:14 scale autonomous vehicle comprising four modular platforms, each housing distinct components of the vehicle. It offers mechatronic redundancy via comprehensive autonomy features as well as an independent 4WD4WS configuration (refer Fig. 3).
@@
Sensor Suite: Nigel hosts a comprehensive sensor suite comprising throttle/steering feedbacks and 1920 CPR incremental encoders for all 4 wheels, a microphone, a 3-axis indoor-positioning system (IPS) using retroreflective/fiducial markers, a 9-axis inertial measurement unit (IMU) with raw/filtered measurements, multiple RGB/RGB-D/stereo camera(s) in the front/rear, and a ${}^{\circ}$ FOV planar LIDAR.
Computation, Communication and Software: Nigel adopts Jetson Orin Nano Developer Kit for most of its high-level computation (autonomy algorithms) and communication (V2V and V2I). Additionally, it also hosts an Arduino Mega (running the vehicle firmware) for acquiring and filtering raw sensor data, and controlling the actuators/lights/indicators.
Actuators: Nigel is actuated using four 6V 160 RPM rated 120:1 DC geared motors to drive its wheels, and four 9.4 kgf.cm servo motors to steer them; the steering actuators are saturated at $\pm$ ${}^{\circ}$ w.r.t. zero-steer value. All the actuators are operated at 5V, which translates to a maximum speed of $\sim$130 RPM for driving and $\sim$0.19 ${}^{\circ}$ for steering. The steering actuators are positioned directly above the respective tire contact patch, which enables zero camber gain or scrubbing while steering the wheels and keeps the actuator effort stable.
@@
Lastly, it is worth mentioning that Nigel supports modular open system architecture (MOSA) standards. This allows the users to adopt and adapt the vehicle for their custom use cases by adding, removing or replacing selective components, or extensively modifying various aspects of the vehicle design.
@@ III Dynamics Modeling and Analysis
Considering the notations presented in Fig. 4, and following the seminal works [38, 39], we can deduce the non-linear yaw-plane vehicle dynamics model for an independent 4WD4WS configuration as depicted in Eq. 1a-1c.
$\mathbf{\sum F_{x}:\>}$ $m\left(\dot{v}\cos(\beta)-v\dot{\beta}\sin(\beta)-\dot{\psi}v\sin(\beta)\right)=\left[\frac{\tau_{RL}}{r_{RL}}\cos(\delta_{RL})+\frac{\tau_{RR}}{r_{RR}}\cos(\delta_{RR})\right.$ (1a)
$+\left.\frac{\tau_{FL}}{r_{FL}}\cos(\delta_{FL})+\frac{\tau_{FR}}{r_{FR}}\cos(\delta_{FR})\right]-\left\{\mu_{RL}C_{RL}\left[\delta_{RL}-\right.\right.$
$\left.\left.\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\sin(\delta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\right.$
$\sin(\delta_{RR})+\left.\mu_{FL}C_{FL}\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\sin(\delta_{FL})+\right.$
$\left.\mu_{FR}C_{FR}\left[\delta_{FR}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\sin(\delta_{FR})\right\}-F_{hw}$
$\mathbf{\sum F_{y}:\>}$ $m\left(\dot{v}\sin(\beta)+v\dot{\beta}\cos(\beta)+\dot{\psi}v\cos(\beta)\right)=\left[\frac{\tau_{RL}}{r_{RL}}\sin(\delta_{RL})+\frac{\tau_{RR}}{r_{RR}}\sin(\delta_{RR})\right.$ (1b)
$+\left.\frac{\tau_{FL}}{r_{FL}}\sin(\delta_{FL})+\frac{\tau_{FR}}{r_{FR}}\sin(\delta_{FR})\right]+\left\{\mu_{RL}C_{RL}\left[\delta_{RL}-\right.\right.$
$\left.\left.\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\right.$
$\cos(\delta_{RR})+\left.\mu_{FL}C_{FL}\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FL})+\right.$
$\left.\mu_{FR}C_{FR}\left[\delta_{FR}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FR})\right\}+F_{sw}$
$\mathbf{\sum M_{z}:\>}$ $I_{z}\ddot{\psi}=l_{f}\left\{-\frac{\tau_{FL}}{r_{FL}}\sin(\delta_{FL}-\theta_{FL})+\frac{\tau_{FR}}{r_{FR}}\sin(\delta_{FR}+\theta_{FR})+\mu_{FL}C_{FL}\right.$ (1c)
$\left.\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FL}-\theta_{FL})+\mu_{FR}C_{FR}\left[\delta_{FR}-\right.\right.$
$\left.\left.\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FR}+\theta_{FR})\right\}-l_{r}\left\{\frac{\tau_{RL}}{r_{RL}}\sin(\delta_{RL}+\theta_{RL})-\right.$
$\left.\frac{\tau_{RR}}{r_{RR}}\sin(\theta_{RR}-\delta_{RR})+\mu_{RL}C_{RL}\left[\delta_{RL}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{RL}+\right.$
$\left.\theta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\theta_{RR}-\delta_{RR})\right\}$
$+\left(\frac{l_{f}-l_{r}}{2}\right)F_{sw}$
where, $F_{y_{i}}=\mu_{i}C_{i}\alpha_{i}$ and $F_{d_{i}}=\begin{cases}\tau_{i}/r_{i};&\text{if }\tau_{i}/r_{i}\leq\mu_{i}N_{i}\\ \mu_{i}N_{i};&\text{otherwise}\end{cases}$ are tire forces and drive forces for $i$-th wheel, respectively. The tire slip angle $\alpha_{i}=\delta_{i}-\beta_{i}$, where $\beta_{i}=\tan^{-1}\left(\frac{\dot{y}\pm\dot{\psi}l_{f/r}}{\dot{x}\pm\dot{\psi}l_{t}}\right)$.
This model can be linearized by applying the small-angle approximation so that $\sin(\odot)\approx\odot$, $\cos(\odot)\approx 1$ and $\tan^{-1}(\odot)\approx\odot$. Additionally, we assume that the vehicle is driving at a constant velocity, implying $\dot{v}\approx 0\Rightarrow\tau_{i}\approx 0$. Finally, it is also assumed that $\dot{x}\gg\dot{\psi}l_{t}\Rightarrow\dot{x}\pm\dot{\psi}l_{t}\approx\dot{x}$. Also note that $F_{sw}$ is hereafter referred to as $F_{w}$ for simplicity, since $F_{hw}$ does not affect the linearized model.
$\begin{bmatrix}\dot{\beta}\\ \ddot{\psi}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{bmatrix}\begin{bmatrix}\beta\\ \dot{\psi}\end{bmatrix}+\begin{bmatrix}b_{11}&b_{12}&b_{13}&b_{14}\\ b_{21}&b_{22}&b_{23}&b_{24}\end{bmatrix}\begin{bmatrix}\delta_{FL}\\ \delta_{FR}\\ \delta_{RL}\\ \delta_{RR}\end{bmatrix}+\begin{bmatrix}d1\\ d2\end{bmatrix}F_{w}$ (2)
where,
@@
$a_{12}=\left\{\frac{1}{mv^{2}}\left[l_{r}\left(\mu_{RL}C_{RL}+\mu_{RR}C_{RR}\right)-l_{f}\left(\mu_{FL}C_{FL}+\mu_{FR}C_{FR}\right)\right]\right\}-1$

### v3

@@ Abstract
Simulation to reality (sim2real) transfer from a dynamics and controls perspective usually involves re-tuning or adapting the designed algorithms to suit real-world operating conditions, which often violates the performance guarantees established originally. This work presents a generalizable framework for achieving reliable sim2real transfer of autonomy-oriented control systems using multi-model multi-objective robust optimal control synthesis, which lends well to uncertainty handling and disturbance rejection with theoretical guarantees. Particularly, this work is centered around a novel actuation-redundant scaled autonomous vehicle called Nigel, with independent all-wheel drive and independent all-wheel steering architecture, whose enhanced configuration space bodes well for robust control applications. To this end, we present the mechatronic design, dynamics modeling, parameter identification, and robust stabilizing as well as tracking control of Nigel using the proposed framework, with exhaustive experimentation and benchmarking in simulation as well as real-world settings.
@@
Autonomous vehicles, over-actuated systems, mechatronic design, robust optimal control, sim2real transfer.
@@ I Introduction
Modern day automotive systems exploit a combination of mechanical electrical, electronic, networking, and software sub-systems to enhance performance via a hierarchical suite of autonomy-oriented control realizations. While earlier autonomy developers may have enjoyed the freedom of primarily focusing on core software development, the present context demands a paradigm shift towards a synergistic hardware-software co-design approach, harmonizing with the principles of mechatronics engineering [1]. In particular, the demand for increased maneuverability, enhanced control configuration space, and improved tolerance against faults motivates the pursuit of unconventional vehicle designs. However, with the advent of novel design architectures, advanced control strategies [2] are required to fully exploit the added capabilities. Moreover, the devised control systems need to successfully transition the sim2real gap by guaranteeing robust performance against real-world uncertainties and disturbances. This further calls for capable cyber-physical deployment platforms of varying scales. Previous works have tried to address some of these aspects in isolation, as we will review below.
Figure 1: Nigel: a mechatronically redundant and reconfigurable 1:14 scale autonomous vehicle tracking “figure-8” maneuver, with the inset depicting disturbance $F_{w}$ and frictional uncertainties $\mu$ during sim2real transfer.
@@
Platform
Cost Sensing Modalities Computational Resources
@@
2D Camera
3D Camera
@@
MIT Racecar [3] 1:10 \$2,600 Jetson TX2 VESC AS
AutoRally [4] 1:5 \$23,300 Custom Teensy LC AS
F1TENTH [5] 1:10 \$3,260 Jetson TX2 VESC 6MkV AS
DSV [6] 1:10 \$1,000 ODROID-XU4 Arduino Mega AS
MuSHR [7] 1:10 \$930 Jetson Nano Turnigy SK8-ESC AS
BARC [8] 1:10 \$1,030 ODROID-XU4 Arduino Nano AS
ORCA [9] 1:43 \$960 None ARM Cortex M4 AS
HyphaROS [10] 1:10 \$600 ODROID-XU4 ESC TBLE-02S AS
Donkey Car [11] 1:16 \$370 Raspberry Pi ESC AS
QCar [12] 1:10 \$20,000 Jetson TX2 Proprietary AS
DeepRacer [13] 1:18 \$400 Proprietary Proprietary AS
Duckiebot [14] N/A \$450 Raspberry Pi None DD
TurtleBot3 [15] N/A \$590 Raspberry Pi OpenCR DD
All costs are ceiled to the nearest \$10. Actuation mechanisms comprise Ackermann-steered (AS), differential-drive (DD), and 4-wheel-drive 4-wheel-steer (4WD4WS) configurations.
@@
In terms of developing scaled autonomous vehicles, academic institutions have contributed platforms like [3, 4, 5, 7, 9, 6, 8]. Community-driven projects such as [10, 11] have also emerged, which are often tailored for specific applications. Moreover, commercial products like [12, 13] have also entered the market, but their closed-source nature and prohibitive costs limit their accessibility to the broader community. Additionally, scaled robot platforms like [15, 14] remain valuable for teaching fundamental autonomy concepts. However, none of the aforementioned platforms (refer Table I) contributes primarily towards novel vehicular configuration or architecture. Other works such as [16, 17] have recently prototyped over-actuated scaled vehicles for validating their control algorithms, but none of them are open-sourced. Additionally, these platforms lack comprehensive autonomy features with only the latter offering truly independent driving and steering capability, although at a much larger scale (1:5) and without extended steering limits. To the best of authors’ knowledge, Nigel (refer Fig. 2) is the first open-source [^1] GitHub: https://github.com/AutoDRIVE-Ecosystem mechatronically redundant [18] autonomous vehicle platform offering comprehensive autonomy features as well as independent all-wheel driving and independent all-wheel steering (i.e., independent 4WD4WS) configuration with extended (±$\pm 90^{\circ}$) steering angles, within a small footprint of 1:14 scale. Additionally, Nigel is a part of the larger AutoDRIVE Ecosystem [^2] Website: https://autodrive-ecosystem.github.io [19, 20], which also offers a high-fidelity digital-twin simulation platform [21, 22], as well as flexible application programming interfaces (APIs) to develop low and high-level autonomy algorithms.
@@
In the context of developing control strategies for vehicles with unconventional architectures, prior works have applied techniques such as inverse dynamics control [23], coordinated motion control [24], adaptive steering control [25] and model predictive control [26, 27]. Although some of these works discuss the adaptability and robustness of the designed controllers in simulation, they fail to guarantee similar performance in real-world conditions. Other recent works such as [28] and [29] have applied cascaded feedforward control and robust $H_{\infty}$ control, respectively, to handle environmental uncertainties and disturbances using over-actuated vehicle architectures. Similar robust control techniques have also been applied to automotive systems for yaw-plane [30, 31, 32], roll-plane [33] or vertical dynamics stabilization [34], emergency braking [35], trajectory tracking [36], etc. However, mixed $H_{2}$-$H_{\infty}$ robust optimal control has not been explored in the automotive field, especially for 4WD4WS autonomous vehicles, and our work seeks to fill in this void.
@@
Specifically, our work seeks to develop a generalizable framework for reliably bridging the sim2real gap by applying the rich theory of multi-model multi-objective control, which seeks the Pareto-optimal solution of $H_{2}$ and $H_{\infty}$ performance tradeoff with $D$-stability guarantee. It is to be noted that, instead of bridging the sim2real gap by increasing simulation fidelity [37], this work intentionally widens the reality gap by using a linearized dynamics model to synthesize the robust optimal controller, which is hypothesized to bridge the dynamics gap during real-world deployment. To this end, we benchmark and validate robust stabilizing, as well as steady-state tracking control of Nigel with seamless sim2real transfer.
To recapitulate, our contributions are itemized below:
@@
  • Completely open-source design architecture of a novel 1:14 scale independent 4WD4WS autonomous vehicle with extended steering limits is presented.
@@
  • Non-linear dynamics model of such a novel vehicle architecture is derived and transformed into a linear uncertain system model with reasonable parameter identification.
@@
  • Robust sim2real control framework is established and validated through exhaustive experimentation and benchmarking in simulation as well as real-world settings.
@@ II Vehicle Design Architecture
Chassis: Nigel is a 1:14 scale autonomous vehicle comprising four modular platforms, each housing distinct components of the vehicle.
@@
Sensor Suite: Nigel hosts a comprehensive sensor suite comprising throttle/steering feedbacks and 1920 CPR incremental encoders for all 4 wheels, a microphone, a 3-axis indoor-positioning system (IPS) using retroreflective/fiducial markers, a 9-axis inertial measurement unit (IMU) with raw/filtered measurements, multiple RGB/RGB-D/stereo camera(s) in the front/rear, and a 360 FOV planar LIDAR.
Computation, Communication and Software: Nigel adopts Jetson Orin Nano Developer Kit for most of its high-level autonomy algorithms and V2X communication and hosts Arduino Mega for acquiring and filtering raw sensor data, and controlling the actuators/lights/indicators.
Actuators: Nigel is actuated using four 6V 160 RPM rated 120:1 DC geared motors to drive its wheels, and four 9.4 kgf.cm servo motors to steer them; the steering actuators are saturated at $\pm$ 90 w.r.t. zero-steer value (see Fig. 3). All the actuators are operated at 5V, which translates to a maximum speed of $\sim$130 RPM for driving and $\sim$0.19 s/60 for steering. The steering actuators are positioned directly above the respective tire contact patch, which enables zero camber gain and keeps the actuator effort stable.
Figure 2: Various components and sub-systems of Nigel. Dashed lines indicate occluded or hidden components. A video highlighting key features of Nigel is available at https://youtu.be/UVIShZuZmpg
@@ III Dynamics Modeling and Analysis
Considering the notations presented in Fig. 4, and extending the seminal works [38, 39], we can deduce the non-linear yaw-plane vehicle dynamics model for an independent 4WD4WS configuration as given in Eq. 1a-1c.
$\begin{split}\mathbf{\sum F_{x}:\>}m\left(\dot{v}\cos(\beta)-v\dot{\beta}\sin(\beta)-\dot{\psi}v\sin(\beta)\right)=\left[\frac{\tau_{RL}}{r_{RL}}\cos(\delta_{RL})\right.\\ \left.+\frac{\tau_{RR}}{r_{RR}}\cos(\delta_{RR})+\frac{\tau_{FL}}{r_{FL}}\cos(\delta_{FL})+\frac{\tau_{FR}}{r_{FR}}\cos(\delta_{FR})\right]-\left\{\mu_{RL}C_{RL}\right.\\ \left.\left[\delta_{RL}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\sin(\delta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\right.\right.\\ \left.\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\sin(\delta_{RR})+\left.\mu_{FL}C_{FL}\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\right.\\ \left.\sin(\delta_{FL})+\mu_{FR}C_{FR}\left[\delta_{FR}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\sin(\delta_{FR})\right\}-F_{hw}\end{split}$ (1a)
$\begin{split}\mathbf{\sum F_{y}:\>}m\left(\dot{v}\sin(\beta)+v\dot{\beta}\cos(\beta)-\dot{\psi}v\cos(\beta)\right)=\left[\frac{\tau_{RL}}{r_{RL}}\sin(\delta_{RL})\right.\\ \left.+\frac{\tau_{RR}}{r_{RR}}\sin(\delta_{RR})+\frac{\tau_{FL}}{r_{FL}}\sin(\delta_{FL})+\frac{\tau_{FR}}{r_{FR}}\sin(\delta_{FR})\right]+\left\{\mu_{RL}C_{RL}\right.\\ \left.\left[\delta_{RL}-\tan^{-1}\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\right.\right.\\ \left.\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\delta_{RR})+\left.\mu_{FL}C_{FL}\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\right.\\ \left.\cos(\delta_{FL})+\mu_{FR}C_{FR}\left[\delta_{FR}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FR})\right\}+F_{sw}\end{split}$ (1b)
$\begin{split}\mathbf{\sum M_{z}:\>}I_{z}\ddot{\psi}=l_{f}\left\{-\frac{\tau_{FL}}{r_{FL}}\sin(\delta_{FL}-\theta_{FL})+\frac{\tau_{FR}}{r_{FR}}\sin(\delta_{FR}+\theta_{FR})\right.\\ \left.+\mu_{FL}C_{FL}\left[\delta_{FL}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FL}-\theta_{FL})+\mu_{FR}C_{FR}\right.\\ \left.\left[\delta_{FR}-\tan^{-1}\left(\frac{\dot{y}+\dot{\psi}l_{f}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\delta_{FR}+\theta_{FR})\right\}-l_{r}\left\{\frac{\tau_{RL}}{r_{RL}}\sin(\delta_{RL}\right.\\ \left.+\theta_{RL})-\frac{\tau_{RR}}{r_{RR}}\sin(\theta_{RR}-\delta_{RR})+\mu_{RL}C_{RL}\left[\delta_{RL}-\tan^{-1}\right.\right.\\ \left.\left.\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}-\dot{\psi}l_{t}}\right)\right]\cos(\delta_{RL}+\theta_{RL})+\mu_{RR}C_{RR}\left[\delta_{RR}-\tan^{-1}\right.\right.\\ \left.\left.\left(\frac{\dot{y}-\dot{\psi}l_{r}}{\dot{x}+\dot{\psi}l_{t}}\right)\right]\cos(\theta_{RR}-\delta_{RR})\right\}+\left(\frac{l_{f}-l_{r}}{2}\right)F_{sw}\end{split}$ (1c)
Here, $m$ is the vehicle mass, $v$ is the vehicle velocity, $\beta=\tan^{-1}\left(\frac{\dot{y}\pm\dot{\psi}l_{f/r}}{\dot{x}\pm\dot{\psi}l_{t}}\right)$ is the vehicle slip angle, $\psi$ is the vehicle yaw angle, $\tau_{i}$ are the drive torques, $r_{i}$ are the wheel radii, $\delta_{i}$ are the steering angles, $\mu_{i}$ are the frictional coefficients for each road-tire interconnect, $C_{i}$ are the tire stiffness values, $\alpha_{i}=\delta_{i}-\beta$ are the tire slip angles, $l_{f}$ and $l_{r}$ are the front and rear center-of-gravity (CG) offsets, $I_{z}$ is the vehicle yaw moment of inertia, $\theta_{i}$ are the angular wheel locations w.r.t. CG, while $F_{hw}$ and $F_{sw}$ are respectively the head and side winds. Note that $F_{y_{i}}=\mu_{i}C_{i}\alpha_{i}$ and $F_{d_{i}}=\begin{cases}\tau_{i}/r_{i};&\text{if }\tau_{i}/r_{i}\leq\mu_{i}N_{i}\\ \mu_{i}N_{i};&\text{otherwise}\end{cases}$ are tire forces and drive forces for $i$-th wheel, respectively.
@@
The non-linear model in Eq. 1a-1c can be linearized by applying the small-angle approximation so that $\sin(\odot)\approx\odot$, $\cos(\odot)\approx 1$ and $\tan^{-1}(\odot)\approx\odot$. Additionally, we assume that the vehicle is driving at a constant velocity, implying $\dot{v}\approx 0\Rightarrow\tau_{i}\approx 0$. Finally, it is also assumed that $\dot{x}\gg\dot{\psi}l_{t}\Rightarrow\dot{x}\pm\dot{\psi}l_{t}\approx\dot{x}$. Also note that $F_{sw}$ is hereafter referred to as $F_{w}$ for simplicity, since $F_{hw}$ does not affect the linearized model.
$\begin{bmatrix}\dot{\beta}\\ \ddot{\psi}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{bmatrix}\begin{bmatrix}\beta\\ \dot{\psi}\end{bmatrix}+\begin{bmatrix}b_{11}&b_{12}&b_{13}&b_{14}\\ b_{21}&b_{22}&b_{23}&b_{24}\end{bmatrix}\begin{bmatrix}\delta_{FL}\\ \delta_{FR}\\ \delta_{RL}\\ \delta_{RR}\end{bmatrix}+\begin{bmatrix}d_{1}\\ d_{2}\end{bmatrix}F_{w}$ (2)
where
@@
$a_{12}=$$\left\{\frac{1}{mv^{2}}\left[l_{r}\left(\mu_{RL}C_{RL}+\mu_{RR}C_{RR}\right)-l_{f}\left(\mu_{FL}C_{FL}\right.\right.\right.$
$\left.\left.\left.+\mu_{FR}C_{FR}\right)\right]\right\}-1$